# Engineering Carbon Credits Towards A Responsible FinTech Era: The Practices, Implications, and Future

### v1

@@ 3 For FinTech: The Impact of Disclosing Carbon Emissions on organisations
In addition to exploring the impact of carbon emissions on corporate market value, our research also found that carbon emission information disclosure and actual emission levels significantly influence corporate loan interest rates. The study indicates that organisations that voluntarily disclose carbon emission information typically enjoy lower loan spreads, particularly in organisations with lower information transparency (such as those with incomplete financial information or insufficient credit ratings). For these organisations, voluntary disclosure of carbon emission information can significantly reduce loan spreads, thereby lowering borrowing costs. Furthermore, the study reveals a positive correlation between actual corporate carbon emission levels, particularly direct carbon emissions (i.e., Scope 1 emissions), and loan spreads. This means that higher carbon emission levels lead to higher loan spreads. Financial markets, especially banks, consider environmental performance, particularly carbon emission levels, when assessing corporate credit risk. High carbon emissions are viewed as potential environmental risks, leading to an increase in loan spreads [9].
@@ 4 For Engineering: Factors Influencing Carbon Credit Prices
However, solely relying on proactive management and disclosure is far from sufficient. To further optimize carbon management strategies, organisations need to purchase carbon credits when prices are low to reduce operational costs and maximize the economic benefits of carbon management. Research has found that for high-emission industries, carbon management costs will constitute a significant portion of expenses. Moreover, if carbon management costs are too high, the company’s competitive position will be weakened [19]. Therefore, understanding and analyzing the various factors that influence carbon credit prices is crucial. Through qualitative analysis of these factors, organisations can preliminarily predict carbon price trends, allowing them to more accurately determine the optimal timing for purchasing credits. Additionally, these factors can serve as input variables for carbon price prediction algorithms, enhancing the accuracy and applicability of predictions.
@@
Coal Rising coal prices lead industries to seek alternative energy sources with fewer carbon emissions, like oil or renewable energy, reducing the demand for carbon credits and thus reducing carbon prices. Econometric Model [24]
Static and Dynamic Panel Models [25]
LASSO and SVAR Models [26]
Multiscale Analysis Model [27]
Vector Error Correction Model (VEC) and GARCH [28]
OLS, IV, VAR model [29]
Econometric Model [30]
SVAR model [31]
@@
Rising coal prices are typically associated with increased demand for traditional energy sources, leading to higher carbon emissions and increased carbon prices. BP Neural Network Model [32]
SEM and BN [33]
Feasible Quasi Generalized Least Squares (FQGLS) [34]
@@
Crude Oil Oil price increases are closely related to overall economic activity and have a strong positive impact on carbon prices. Econometric Model [24]
BP Neural Network Model [32]
SV-TVP-VAR Model [35]
SEM and BN [33]
MEEMD and Fast Fourier Transform [36]
TVP-VAR Model [37]
Feasible Quasi Generalized Least Squares (FQGLS) [34]
SVAR model [31]
@@
Increases in crude oil prices raise production costs, leading organisations to reduce supply, which lowers carbon emissions and reduces the demand and price of carbon emission futures. Vector Error Correction (VEC) Model [38]
SV-TVP-VAR Model [35]
@@
Macroeconomics Growth in macroeconomic activity typically drive increases in industrial production and raw material consumption, leading to higher demand for carbon emissions and pushing carbon prices up. Econometric Model [24]
Vector Error Correction (VEC) Model [38]
BP Neural Network Model [32]
MEMD, VAR, VEC Models [39]
EEMD (Ensemble Empirical Mode Decomposition) [40]
Vector Error Correction Model (VEC) and GARCH [28]
SVAR model [31]
@@
Similar Products An increase in similar products’ prices often leads to a rise in carbon credit prices, indicating a substitution relationship between the two. Econometric Model [24]
SV-TVP-VAR Model [35]
ARMA-GARCH [41]
@@
Stock Market Indices Stock market indices have a significant positive impact on the carbon market SEM and BN [33]
Static and Dynamic Panel Models [25]
LASSO and SVAR Models [26]
ARIMA-GARCH Model, Random Forest Algorithm [42]
Multiscale Analysis Model [27]
@@
Euro Index The Euro Index had a negative impact on carbon prices before COVID-19, mainly due to uncertainties like Brexit MEEMD and Fast Fourier Transform [36]
@@
Utility Index The utility index positively impacts carbon prices, reflecting increased power plant activity and associated carbon emissions. MEEMD and Fast Fourier Transform [36]
@@
Interbank lending rates Interbank lending rates negatively correlate with carbon prices; rising rates weaken market liquidity and reduce demand for carbon allowances, leading to lower carbon prices. Bai-Perron Structural Break Test and Newey-West Regression Model [43]
@@
Exchange rate Exchange rates negatively impact carbon prices, particularly in the short and medium term. DCC-GARCH, TVP-VAR and GA-BP [44]
MEMD, VAR, VEC Models [39]
@@
COVID-19 Pandemic The COVID-19 pandemic significantly reduced carbon prices. Bai-Perron Structural Break Test and Newey-West Regression Model [43]
EEMD (Ensemble Empirical Mode Decomposition) [40]
@@
Clean Energy Rising clean energy prices will drive up carbon prices, while the development of clean energy may sometimes suppress carbon prices. BP Neural Network Model [32]
SEM and BN [33]
MEEMD and Fast Fourier Transform [36]
Static and Dynamic Panel Models [25]
TVP-VAR Model [37]
OLS, IV, VAR model [29]
Feasible Quasi Generalized Least Squares (FQGLS) [34]
Econometric Model [30]
SVAR model [31]
@@
Electricity price The impact of electricity prices positively influence the price of carbon credits. TVP- VAR Model [37]
Multiscale Analysis Model [27]
OLS, IV, VAR model [29]
Econometric Model [30]
@@
End product prices The rise in prices of end products such as minerals, steel, and paper can drive up carbon credit prices, as the increase in end product prices leads to higher production, thereby increasing the demand for carbon emissions. OLS, IV, VAR model [29]
@@
Air quality Environmental factors like air quality significantly impact carbon prices; higher air quality index (AQI) levels are associated with rising carbon prices. ARMA-GARCH [41]
VAR-VEC model, GARCH model [45]
Vector Error Correction Model (VEC) and GARCH [28]
@@
Temperature Temperature has a significant impact on carbon allowance trading volume. Extreme weather conditions, such as very cold or hot temperatures, increase energy demand, thereby affecting the carbon allowance market and driving up carbon prices. Econometric Model [30]
Autoregressive Distributed Lag Model (ARDL) [46]
Structural Break Test, Autoregressive Distributed Lag Model (ARDL) [47]
@@
Supply factor The supply of carbon credits also affects carbon prices; when the supply is large, carbon prices decrease. Econometric Model [30]
@@
Degree of Carbon Regulation Short-term negative impact of carbon regulation policies on carbon prices is greater than long-term effects, but carbon prices become more sensitive to policy information as market mechanisms improve. SV-TVP-VAR Model [35]
@@
Interest Rates Interest rates have a negative short-term impact on carbon prices SV-TVP-VAR Model [35]
@@
Interest rates have a positive impact on carbon prices, especially in the medium to long term, as higher interest rates significantly reduce carbon prices by increasing production costs. DCC-GARCH, TVP-VAR and GA-BP [44]
@@
Gold Prices Gold prices positively affect carbon prices, indicating a relationship between increased gold mining activities and higher carbon emissions. MEEMD and Fast Fourier Transform [36]
@@
Population growth rate There is a positive correlation between population growth rate and carbon prices. A higher population growth rate will increase future carbon prices, as population growth leads to higher carbon emissions, necessitating higher carbon prices to address climate change. EZ Climate Model and Uncertainty Analysis [48]
@@
Growth rate of per capita output The growth rate of per capita output has a greater impact on carbon prices. A higher per capita output growth rate will significantly increase future carbon prices. EZ Climate Model and Uncertainty Analysis [48]
@@
Energy prices play a crucial role in determining carbon credit prices. Increases in coal and crude oil prices typically lead to higher carbon emissions, which in turn drive up carbon credit prices. However, as coal prices rise, industrial sectors may shift towards alternative energy sources with lower carbon emissions, such as natural gas or renewable energy, which could reduce the demand for carbon credits and subsequently lower prices. The impact of natural gas prices on carbon credits can be twofold. On one hand, higher natural gas prices may elevate carbon credit prices. On the other hand, the development of renewable energy may temper this upward trend [24, 32, 35, 33, 36, 25, 26, 37, 27]. The above conclusion has been confirmed in multiple studies. For example, [26] justifies the negative relationship between coal prices and carbon emissions using the LASSO and SVAR models.
@@
Macroeconomic indicators and stock market performance also significantly influence carbon credit prices. Economic growth often acorganisations increased industrial production and energy consumption, thereby raising the demand for carbon emissions and driving up carbon credit prices. Moreover, stock market indices, such as the Euro Stoxx 50 and S&P 500, play a critical role in the carbon market, particularly during periods of economic prosperity. However, in certain markets, such as China, the relationship between stock market performance and carbon prices may be negative, indicating that market maturity and investor expectations influence price fluctuations [24, 38, 32, 33, 25, 26, 42, 39, 27, 40]. Numerous studies have also validated the relationship between economic indicators and carbon prices. For example, [38] justifies the positive relationship between macroeconomic development and carbon prices using the Vector Error Correction (VEC) Model.
@@
Regulatory policies have a significant impact on the carbon market. In the short term, stringent carbon regulation policies may increase compliance costs for organisations, thereby depressing carbon prices. However, as market mechanisms improve, the sensitivity of carbon prices to policy information may increase, potentially leading to positive impacts in the long run. Additionally, the price linkage between international carbon markets is another key consideration, particularly the positive impact of European carbon markets on China’s carbon market, which has become more evident since 2020 [35, 25, 43, 37].
@@
Financial factors such as exchange rates and interest rates also have complex effects on carbon credit prices. Exchange rate fluctuations generally have a negative impact on carbon prices in the short and medium term. For example, when the Renminbi depreciates, carbon prices may rise in the short term but decline in the long term. Rising interest rates typically suppress market liquidity, reducing demand for carbon credits and leading to lower prices. However, in the medium to long term, rising interest rates may increase production costs, reducing demand for carbon emissions and thereby pushing up carbon prices [43, 44, 39]. These conclusions have also been validated in several studies. For example, [43] justifies the negative relationship between interbank lending rates and carbon prices using the Bai-Perron Structural Break Test and Newey-West Regression Model.
@@
Market sentiment and significant events, such as the COVID-19 pandemic, have profound effects on carbon credit price volatility. During the pandemic, global economic uncertainty and supply chain disruptions led to a sharp decline in carbon prices. However, the implementation of the European Union’s 750 billion euro green recovery plan helped stabilize the carbon market, leading to a rebound in carbon prices. Additionally, fluctuations in market sentiment may exhibit nonlinear relationships, adding complexity to price predictions. In such cases, nonlinear models like random forests have demonstrated superior accuracy in predicting price changes [36, 43, 44, 41, 24].
@@
Air quality is a significant environmental factor that influences carbon prices. Higher Air Quality Index (AQI) levels are typically associated with rising carbon prices, as poorer air quality often indicates increased carbon emissions, thereby driving up carbon prices. Several studies have confirmed this relationship [41, 45, 28]. For instance, [41] validated this positive correlation using the ARMA-GARCH model.
@@
Temperature has a notable impact on carbon allowance trading volumes. Extreme weather conditions, such as very cold or hot temperatures, increase energy demand, which in turn affects the carbon allowance market and drives up carbon prices. Multiple studies have validated this effect using various econometric models [30, 46, 47].
@@
The supply of carbon credits also affects carbon prices. When the supply of carbon credits is large, prices tend to decrease due to market oversupply. [30] demonstrated the inverse relationship between supply levels and carbon prices using an econometric model.
@@
Gold prices have a positive effect on carbon prices, as increased gold mining activities are typically associated with higher carbon emissions, which raises the demand for carbon credits. [36] used the MEEMD and Fast Fourier Transform models to demonstrate the positive relationship between gold prices and carbon prices.
@@
There is a positive correlation between the population growth rate and carbon prices. A higher population growth rate is expected to increase future carbon prices, as population growth is often accompanied by higher carbon emissions. [48] used the EZ Climate Model and Uncertainty Analysis to demonstrate the impact of population growth on carbon prices.
@@
The growth rate of per capita output significantly influences carbon prices. Higher per capita output growth rates will lead to substantial increases in future carbon prices, as this growth is typically accompanied by increased industrial activity and energy consumption. [48] validated the influence of per capita output growth on carbon prices using the EZ Climate Model and Uncertainty Analysis.
@@
In addition to the factors mentioned above, other influences have also been shown to affect carbon prices. For example, [41] justifies the negative relationship between the prices of similar products using the ARMA-GARCH model. Electricity prices have also been confirmed by numerous studies to have a positive impact on carbon prices [37, 27, 29, 30]. For example, [30] used an Econometric Model to validate this effect.
@@
European Union Allowance (EUA) Black Energy (Coal, Oil) and Clean Energy (Natural Gas, Solar) BP Neural Network Model [32]
Gold Price, Euro Index and Utility Index MEMD and FFT [36]
Stock Market Indices (SP 500, CAC 40, DAX) and Carbon Emission Market (EUA, CER) SEM and BN [33]
Market Stability Reserve (MSR) N-A MEMD, VAR, VEC Models [39]
Energy Market and Stock Market LASSO and MSVAR Models [26]
@@
COVID-19 Pandemic Bai-Perron Structural Break Test and Newey-West Regression Model [43]
@@
China’s Carbon Market Degree of Carbon Regulation, Crude Oil Price, EUA Futures Price and Interest Rates SV-TVP-VAR Model [35]
Exchange Rate and Coal Price DCC-GARCH, TVP-VAR and GA-BP [44]
@@
Industrial Production, Stock Market Index and Renewable Energy Generation Static and Dynamic Panel Models [25]
@@
Industrial Structure, Policy Factors, Environmental Factors and External Market ARMA-GARCH [41]
@@
Energy Prices, Economic Indicators, Macroeconomic Variables and Market Sentiment ARIMA(1,1,1)-GARCH(1,1) Model, Random Forest Algorithm [42]
Significant Short-term Events EEMD [40]
@@
Table 2 summarizes the key factors influencing the European Union Allowance (EUA) market and China’s carbon market, along with the relevant research models and methodologies. This table clearly illustrates the differences in influencing factors across these markets.
@@
In the EUA market, carbon prices are affected by energy prices, financial market indicators, policy factors, and external events. In terms of the energy market, price fluctuations of black energy (such as coal and crude oil) and clean energy (such as natural gas and solar) are major drivers of carbon prices. As traditional energy prices rise, organisations tend to shift toward low-carbon energy, thereby affecting the supply and demand of carbon credits [32]. Regarding financial market indicators, factors such as gold prices, the Euro Index, and the utility index also influence carbon prices [33, 36, 26]. These indicators reflect macroeconomic and financial market changes, and their connection to the energy market has been verified through models such as MEMD and FFT [36]. As for stock market and carbon market indices, stock market indices (such as SP 500, CAC 40, DAX) have a significant positive impact on the carbon market, as they reflect economic growth expectations, which in turn affect energy demand and carbon emissions [33, 26]. Additionally, the Market Stability Reserve (MSR) is a policy mechanism unique to the EUA market, designed to stabilize the market and prevent excessive carbon price fluctuations, with its impact analyzed through VAR and VEC models [39]. The COVID-19 pandemic, as an external event, caused short-term shocks to the carbon market, most notably the sharp decline in carbon prices in 2020, followed by a recovery driven by stimulus plans [43].

### v2

@@ 3 For Finance: The Impact of Disclosing Carbon Emissions on organisations
In addition to exploring the impact of carbon emissions on corporate market value, our research also found that carbon emission information disclosure and actual emission levels significantly influence corporate loan interest rates. The study indicates that organisations that voluntarily disclose carbon emission information typically enjoy lower loan spreads, particularly in organisations with lower information transparency (such as those with incomplete financial information or insufficient credit ratings). For these organisations, voluntary disclosure of carbon emission information can significantly reduce loan spreads, thereby lowering borrowing costs. Furthermore, the study reveals a positive correlation between actual corporate carbon emission levels, particularly direct carbon emissions (i.e., Scope 1 emissions), and loan spreads. This means that higher carbon emission levels lead to higher loan spreads. Financial markets, especially banks, consider environmental performance, particularly carbon emission levels, when assessing corporate credit risk. High carbon emissions are viewed as potential environmental risks, leading to an increase in loan spreads [carbon72].
@@ 4 For Engineering: Factors Influencing Carbon Credit Prices
However, solely relying on proactive management and disclosure is far from sufficient. To further optimize carbon management strategies, organisations need to purchase carbon credits when prices are low to reduce operational costs and maximize the economic benefits of carbon management. Research has found that for high-emission industries, carbon management costs will constitute a significant portion of expenses. Moreover, if carbon management costs are too high, the company’s competitive position will be weakened [carbon81]. Therefore, understanding and analyzing the various factors that influence carbon credit prices is crucial. Through qualitative analysis of these factors, organisations can preliminarily predict carbon price trends, allowing them to more accurately determine the optimal timing for purchasing credits. Additionally, these factors can serve as input variables for carbon price prediction algorithms, enhancing the accuracy and applicability of predictions.
@@
Coal Rising coal prices lead industries to seek alternative energy sources with fewer carbon emissions, like oil or renewable energy, reducing the demand for carbon credits and thus reducing carbon prices. Econometric Model [n1]
Static and Dynamic Panel Models [n7]
LASSO and SVAR Models [n8]
Multiscale Analysis Model [n15]
Vector Error Correction Model (VEC) and GARCH [n17]
OLS, IV, VAR model [n18]
Econometric Model [n20]
SVAR model [n21]
@@
Rising coal prices are typically associated with increased demand for traditional energy sources, leading to higher carbon emissions and increased carbon prices. BP Neural Network Model [n3]
SEM and BN [n5]
Feasible Quasi Generalized Least Squares (FQGLS) [n19]
@@
Crude Oil Oil price increases are closely related to overall economic activity and have a strong positive impact on carbon prices. Econometric Model [n1]
BP Neural Network Model [n3]
SV-TVP-VAR Model [n4]
SEM and BN [n5]
MEEMD and Fast Fourier Transform [n6]
TVP-VAR Model [n10]
Feasible Quasi Generalized Least Squares [n19]
SVAR model [n21]
@@
Increases in crude oil prices raise production costs, leading organisations to reduce supply, which lowers carbon emissions and reduces the demand and price of carbon emission futures. Vector Error Correction Model [n2]
SV-TVP-VAR Model [n4]
@@
Macroeconomics Growth in macroeconomic activity typically drive increases in industrial production and raw material consumption, leading to higher demand for carbon emissions and pushing carbon prices up. Econometric Model [n1]
Vector Error Correction Model [n2]
BP Neural Network Model [n3]
MEMD, VAR, VEC Models [n14]
Ensemble Empirical Mode Decomposition (EEMD) [n16]
Vector Error Correction Model and GARCH [n17]
SVAR model [n21]
@@
Similar Products An increase in similar products’ prices often leads to a rise in carbon credit prices, indicating a substitution relationship between the two. Econometric Model [n1]
SV-TVP-VAR Model [n4]
ARMA-GARCH [n12]
@@
Stock Market Indices Stock market indices have a significant positive impact on the carbon market. SEM and BN [n5]
Static and Dynamic Panel Models [n7]
LASSO and SVAR Models [n8]
ARIMA-GARCH Model, Random Forest Algorithm [n13]
Multiscale Analysis Model [n15]
@@
Euro Index The Euro Index had a negative impact on carbon prices before COVID-19, mainly due to uncertainties like Brexit. MEEMD and Fast Fourier Transform [n6]
@@
Utility Index The utility index positively impacts carbon prices, reflecting increased power plant activity and associated carbon emissions. MEEMD and Fast Fourier Transform [n6]
@@
Interbank lending rates Interbank lending rates negatively correlate with carbon prices; rising rates weaken market liquidity and reduce demand for carbon allowances, leading to lower carbon prices. Bai-Perron Structural Break Test and Newey-West Regression Model [n9]
@@
Exchange rate Exchange rates negatively impact carbon prices, particularly in the short and medium term. DCC-GARCH, TVP-VAR and GA-BP [n11]
MEMD, VAR, VEC Models [n14]
@@
COVID-19 Pandemic The COVID-19 pandemic significantly reduced carbon prices. Bai-Perron Structural Break Test and Newey-West Regression Model [n9]
EEMD [n16]
@@
Clean Energy Rising clean energy prices will drive up carbon prices, while the development of clean energy may sometimes suppress carbon prices. BP Neural Network Model [n3]
SEM and BN [n5]
MEEMD and Fast Fourier Transform [n6]
Static and Dynamic Panel Models [n7]
TVP-VAR Model [n10]
OLS, IV, VAR model [n18]
Feasible Quasi Generalized Least Squares [n19]
Econometric Model [n20]
SVAR model [n21]
@@
Electricity price The impact of electricity prices positively influence the price of carbon credits. TVP-VAR Model [n10]
Multiscale Analysis Model [n15]
OLS, IV, VAR model [n18]
Econometric Model [n20]
@@
End product prices The rise in prices of end products such as minerals, steel, and paper can drive up carbon credit prices, as the increase in end product prices leads to higher production, thereby increasing the demand for carbon emissions. OLS, IV, VAR model [n18]
@@
Air quality Environmental factors like air quality significantly impact carbon prices; higher air quality index (AQI) levels are associated with rising carbon prices. ARMA-GARCH [n12]
VAR-VEC model and GARCH [n22]
Vector Error Correction Model and GARCH [n17]
@@
Temperature Temperature has a significant impact on carbon allowance trading volume. Extreme weather conditions, such as very cold or hot temperatures, increase energy demand, thereby affecting the carbon allowance market and driving up carbon prices. Econometric Model [n20]
Autoregressive Distributed Lag Model (ARDL) [n23]
Structural Break Test [n24]
@@
Supply factor The supply of carbon credits also affects carbon prices; when the supply is large, carbon prices decrease. Econometric Model [n20]
@@
Degree of Carbon Regulation Short-term negative impact of carbon regulation policies on carbon prices is greater than long-term effects, but carbon prices become more sensitive to policy information as market mechanisms improve. SV-TVP-VAR Model [n4]
@@
Interest Rates Interest rates have a negative short-term impact on carbon prices. SV-TVP-VAR Model [n4]
@@
Interest rates have a positive impact on carbon prices, especially in the medium to long term, as higher interest rates significantly reduce carbon prices by increasing production costs. DCC-GARCH, TVP-VAR and GA-BP [n11]
@@
Gold Prices Gold prices positively affect carbon prices, indicating a relationship between increased gold mining activities and higher carbon emissions. MEEMD and Fast Fourier Transform [n6]
@@
Population growth rate There is a positive correlation between population growth rate and carbon prices. A higher population growth rate will increase future carbon prices, as population growth leads to higher carbon emissions. EZ Climate Model and Uncertainty Analysis [n25]
@@
Growth rate of per capita output The growth rate of per capita output has a greater impact on carbon prices. A higher per capita output growth rate will significantly increase future carbon prices. EZ Climate Model and Uncertainty Analysis [n25]
@@
Energy prices play a crucial role in determining carbon credit prices. Increases in coal and crude oil prices typically lead to higher carbon emissions, which in turn drive up carbon credit prices. However, as coal prices rise, industrial sectors may shift towards alternative energy sources with lower carbon emissions, such as natural gas or renewable energy, which could reduce the demand for carbon credits and subsequently lower prices. The impact of natural gas prices on carbon credits can be twofold too. On one hand, higher natural gas prices may elevate carbon credit prices. On the other hand, the development of renewable energy may temper this upward trend [n1, n3, n4, n5, n6, n7, n8, n10, n15].
@@
Macroeconomic indicators and stock market performance also significantly influence carbon credit prices. Economic growth often acorganisations increased industrial production and energy consumption, thereby raising the demand for carbon emissions and driving up carbon credit prices. Moreover, stock market indices, such as the Euro Stoxx 50 and S&P 500, play a critical role in the carbon market, particularly during periods of economic prosperity. However, in certain markets, such as China, the relationship between stock market performance and carbon prices may be negative, indicating that market maturity and investor expectations influence price fluctuations [n1, n2, n3, n5, n7, n8, n13, n14, n15, n16].
@@
Regulatory policies have a significant impact on the carbon market. In the short term, stringent carbon regulation policies may increase compliance costs for organisations, thereby depressing carbon prices. However, as market mechanisms improve, the sensitivity of carbon prices to policy information may increase, potentially leading to positive impacts in the long run. Additionally, the price linkage between international carbon markets is another key consideration, particularly the positive impact of European carbon markets on China’s carbon market, which has become more evident since 2020 [n4, n7, n9, n10].
@@
Financial factors such as exchange rates and interest rates also have complex effects on carbon credit prices. Exchange rate fluctuations generally have a negative impact on carbon prices in the short and medium term. For example, when the Renminbi depreciates, carbon prices may rise in the short term but decline in the long term. Rising interest rates typically suppress market liquidity, reducing demand for carbon credits and leading to lower prices. However, in the medium to long term, rising interest rates may increase production costs, reducing demand for carbon emissions and thereby pushing up carbon prices [n9, n11, n14]. These conclusions have also been validated in several studies. For example, [n9] justifies the negative relationship between interbank lending rates and carbon prices using the Bai-Perron Structural Break Test and Newey-West Regression Model.
@@
Market sentiment and significant events, such as the COVID-19 pandemic, have profound effects on carbon credit price volatility. During the pandemic, global economic uncertainty and supply chain disruptions led to a sharp decline in carbon prices. However, the implementation of the European Union’s 750 billion euro green recovery plan helped stabilize the carbon market, leading to a rebound in carbon prices. Additionally, fluctuations in market sentiment may exhibit nonlinear relationships, adding complexity to price predictions. In such cases, nonlinear models like random forests have demonstrated superior accuracy in predicting price changes [n6, n9, n11, n12, n1].
@@
Air quality is a significant environmental factor that influences carbon prices. Higher Air Quality Index levels are typically associated with rising carbon prices, as poorer air quality often indicates increased carbon emissions, thereby driving up carbon prices. Several studies have confirmed this relationship [n12, n22, n17]. For instance, [n12] validated this positive correlation using the ARMA-GARCH model.
@@
Temperature has a notable impact on carbon allowance trading volumes. Extreme weather conditions, such as very cold or hot temperatures, increase energy demand, which in turn affects the carbon allowance market and drives up carbon prices. Multiple studies have validated this effect using various econometric models [n20, n23, n24].
@@
The supply of carbon credits also affects carbon prices. When the supply of carbon credits is large, prices tend to decrease due to market oversupply. [n20] demonstrated the inverse relationship between supply levels and carbon prices using an econometric model.
@@
Gold prices have a positive effect on carbon prices, as increased gold mining activities are typically associated with higher carbon emissions, which raises the demand for carbon credits. [n6] used the MEEMD and Fast Fourier Transform models to demonstrate the positive relationship between gold prices and carbon prices.
@@
There is a positive correlation between the population growth rate and carbon prices. A higher population growth rate is expected to increase future carbon prices, as population growth is often accompanied by higher carbon emissions. [n25] used the EZ Climate Model and Uncertainty Analysis to demonstrate the impact of population growth on carbon prices.
@@
The growth rate of per capita output significantly influences carbon prices. Higher per capita output growth rates will lead to substantial increases in future carbon prices, as this growth is typically accompanied by increased industrial activity and energy consumption. [n25] validated the influence of per capita output growth on carbon prices using the EZ Climate Model and Uncertainty Analysis.
@@
In addition to the factors mentioned above, other influences have also been shown to affect carbon prices. For example, [n12] justifies the negative relationship between the prices of similar products using the ARMA-GARCH model. Electricity prices have also been confirmed by numerous studies to have a positive impact on carbon prices [n10, n15, n18, n20].
@@
European Union Allowance (EUA) Black Energy (Coal, Oil) and Clean Energy (Natural Gas, Solar) BP Neural Network Model [n3]
Gold Price, Euro Index and Utility Index MEMD and FFT [n6]
Stock Market Indices (SP 500, CAC 40, DAX) and Carbon Emission Market (EUA, CER) SEM and BN [n5]
Market Stability Reserve (MSR) N-A MEMD, VAR, VEC Models [n14]
Energy Market and Stock Market LASSO and MSVAR Models [n8]
@@
COVID-19 Pandemic Bai-Perron Structural Break Test and Newey-West Regression Model [n9]
@@
China’s Carbon Market Degree of Carbon Regulation, Crude Oil Price, EUA Futures Price and Interest Rates SV-TVP-VAR Model [n4]
Exchange Rate and Coal Price DCC-GARCH, TVP-VAR and GA-BP [n11]
@@
Industrial Production, Stock Market Index and Renewable Energy Generation Static and Dynamic Panel Models [n7]
@@
Industrial Structure, Policy Factors, Environmental Factors and External Market ARMA-GARCH [n12]
@@
Energy Prices, Economic Indicators, Macroeconomic Variables and Market Sentiment ARIMA(1,1,1)-GARCH(1,1) Model, Random Forest Algorithm [n13]
Significant Short-term Events EEMD [n16]
@@
Table 2 summarizes the key factors influencing the European Union Allowance market and China’s carbon market, along with the relevant research models and methodologies. This table clearly illustrates the differences in influencing factors across these markets.
@@
In the EUA market, carbon prices are affected by energy prices, financial market indicators, policy factors, and external events. In terms of the energy market, price fluctuations of black energy (such as coal and crude oil) and clean energy (such as natural gas and solar) are major drivers of carbon prices. As traditional energy prices rise, organisations tend to shift toward low-carbon energy, thereby affecting the supply and demand of carbon credits [n3]. Regarding financial market indicators, factors such as gold prices, the Euro Index, and the utility index also influence carbon prices [n5, n6, n8]. These indicators reflect macroeconomic and financial market changes, and their connection to the energy market has been verified through models such as MEMD and FFT [n6]. As for stock market and carbon market indices, stock market indices (such as SP 500, CAC 40, DAX) have a significant positive impact on the carbon market, as they reflect economic growth expectations, which in turn affect energy demand and carbon emissions [n5, n8]. Additionally, the Market Stability Reserve (MSR) is a policy mechanism unique to the EUA market, designed to stabilize the market and prevent excessive carbon price fluctuations, with its impact analyzed through VAR and VEC models [n14]. The COVID-19 pandemic, as an external event, caused short-term shocks to the carbon market, most notably the sharp decline in carbon prices in 2020, followed by a recovery driven by stimulus plans [n9].